OpenStudy (anonymous):

use a double or half angle formula to solve the equation in the interval [0,2pi) tan(a)+cot(a)=4sin2(a)

OpenStudy (e.mccormick):

Only way to solve is to set to zero. So start with that.

OpenStudy (anonymous):

tan(a)+cot(a)-4sin2(a)=0

OpenStudy (anonymous):

like that?

OpenStudy (e.mccormick):

Yep. Now for the fun part. Seeing which double or half will help the most.

OpenStudy (anonymous):

okay what does that mean?

OpenStudy (e.mccormick):

Well, you generlly want things to factor so that you can set each term to zero. Some things are easier to factor than others.

OpenStudy (anonymous):

Oh it's fixed now.

OpenStudy (e.mccormick):

はい

OpenStudy (anonymous):

so whats next? what?? hah

OpenStudy (e.mccormick):

He picked on me about letting my japanese get overed up by th number in the lower right.... so I had to fix it so the harmony of the aviutar would be maintained.

OpenStudy (anonymous):

hahah ight.

OpenStudy (e.mccormick):

OK, using their equation editor, that is with a power already, right? \[\tan(a)+\cot(a)-4\sin^2(a)=0\]

OpenStudy (anonymous):

man im getting tired haha its later over here on the east coast lol

OpenStudy (e.mccormick):

Hmmm. what about putting tan and cot in terms of sin and cos. Then seeing what pops up that might let a half or double formula (as required) clean it up?

OpenStudy (e.mccormick):

And yah, 3 hour time difference?

OpenStudy (anonymous):

tan is sin/cos and cot is cos/tan?

OpenStudy (e.mccormick):

cot = cos/sin

OpenStudy (anonymous):

woulsnt that cancel out, and it would be -4sin2(a)=0?

OpenStudy (e.mccormick):

Not quite. You can get it to 1/(sin cos) but I think I found a better way. I was playing with it a bit, then decided to dig. Man, I really need to brush up on trig. I have a calc teacher that loves trig...

OpenStudy (anonymous):

okay whats this better way. how bout you just show me how to so it. im about to fall asleep. hah please..

OpenStudy (e.mccormick):

I am going to shortcut this a bit for speed. So you see that we can get to:\[\frac{\sin}{\cos} + \frac{\cos}{\sin}=4\sin^2\]

OpenStudy (anonymous):

okay good.

OpenStudy (e.mccormick):

\[\frac{\sin^2+\cos^2}{\sin \cos}=bla\] that can be used to jump to the double angle of \[\frac{1} {(1/2)\sin(2a))}=bla\]

OpenStudy (e.mccormick):

Ge that step? Well, 2 steps. I pathagoreaned the top and double angled the bottom.

OpenStudy (anonymous):

what does bla stand for? is that the final answer?

OpenStudy (e.mccormick):

Just repeating the same right hand side. I was not changing it, so bla.

OpenStudy (e.mccormick):

Now, can you see how that left hald side would simplify to get rid of the fraction?

OpenStudy (anonymous):

so it equals 4sin2(a)

OpenStudy (e.mccormick):

\[\frac{1}{\frac{\sin(2a)}{2}}=4\sin^2a \rightarrow 2\sin(2a) =4\sin^2a \rightarrow 2\sin(2a) - 4\sin^2a = 0\]

OpenStudy (e.mccormick):

Now that looks more factorable.

OpenStudy (anonymous):

what do i factor out? -2sin(a)?

OpenStudy (e.mccormick):

Hmmm.... I know we are close. Lets see. I may need to back up a step and leave part in the 2sin cos. Lets look at the 4sin^2 now.

OpenStudy (e.mccormick):

With half angle, that is -4(1/2 1-cos(2a))

OpenStudy (e.mccormick):

Yah, so let me put that in there. \[2\sin a \cos a - 2 - 2 \cos(2a) \; or \; 2\sin(2a)-2-2 \cos(2a)\] does anything theremake it better..

OpenStudy (e.mccormick):

If you did not notice, I distributed the 4 into the half angle, which got me the 2-2cos(2a)... but I messed up the sign. It should be +2cos(2a).

OpenStudy (anonymous):

so the answer the sin(2a) -2 +2cos(2a) ?

OpenStudy (e.mccormick):

Well, that is close. But if hey want you to find values of a, it is not an answer. I was aiming for factors to set to 0 and somewhere in there I did not get what I wanted.

OpenStudy (e.mccormick):

OK. I also see this in there: \[2\sin a(\cos a - 2\sin a) = 0\]which is partially factored. But I would need to get the factors of what is inside the parantisis. Which is what I thought was going to be nicer than it ended up.

OpenStudy (e.mccormick):

You understand where I am going with the factoring? This is like algebra. x^2+4x+3 = 0, so you factor it to (x+1)(x+3)=0 and then I can solve each part. That is how you solve this sort of equation, unless you just needed a proof that one side equals the other.

OpenStudy (e.mccormick):

Just to make sure, is that problem \(4sin^2(a)\) or \(4sin^2(2a)\)?

OpenStudy (anonymous):

4sin(2a)

OpenStudy (anonymous):

no squared

OpenStudy (e.mccormick):

OH! Ok. That was my mistake.

OpenStudy (e.mccormick):

Well that makes all the difference!

OpenStudy (anonymous):

can you show me all the work now? cause im all confused

OpenStudy (anonymous):

tan(a) + cot(a) = 4sin(2a) thats the problem

OpenStudy (e.mccormick):

Yah, I saw up above 4sin2(a) and I thought you meant squared. Big goog there on my part for not making sure earlier.

OpenStudy (e.mccormick):

The start was good then. \[\frac{\sin a}{\cos a}+\frac{\cos a}{\sin a}-4\sin(2a)=0 \rightarrow \frac{\sin^2 a + \cos^2 a}{\cos a \sin a}-4\sin(2a)=0\] \[\rightarrow \frac{1}{\cos a \sin a}-4sin(2a)=0\]Wit me to there?

OpenStudy (anonymous):

yeah, is that it?

OpenStudy (e.mccormick):

Not quite. Then I did that step where we went to this (but corrected):\[2\sin(2a)-4\sin(2a)=0\] which simplifies as:\[-2\sin{(2a)}=0\]

OpenStudy (e.mccormick):

So find every part of the interval where \(\sin{(2a)}=0\) and you have the solutions. There should be a small herd of them with that.

OpenStudy (e.mccormick):

On \([0,2\pi)\) you should have something like 4 or 8 points.Lets see, on the unit circle, sine is 0 at 0 and \(\pi\). But this is of (2a) so when 2a=0 or 2a=\(\pi\) you have a zero.

OpenStudy (anonymous):

ok, yeah i got that

OpenStudy (e.mccormick):

I'm at the end of a long work day, and I know you are brain fried at this point, so write it all out and then try it to see if you can get the same thing. If you can; then I did not make any mistakes. But if you can't I may have goofed in there somewhere and you may need someone else to go over it with you.

OpenStudy (e.mccormick):

And the working it out on your part should be after some sleep. Hehe.

OpenStudy (anonymous):

ok sounds good. thank you. whats the exact answer? 0 and π?

OpenStudy (e.mccormick):

Should be a={0,pi/2,pi,3pi/4}

OpenStudy (e.mccormick):

Because of the 2a part.

OpenStudy (anonymous):

thanks man. goodnight

OpenStudy (e.mccormick):

Night.